# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 6. Linearization
Now, it is possible that $\beta(t_{l(T)}^{+})\varphi(t_{l(T)}^{+})\cdot w_{l(T)}\in\mathring{\Phi}$ and $t_{l(T)}^{+}=T$. So, we will estimate the $l(T)$-th term differently. By Proposition 1.15, and recalling (1.6), given $\epsilon>0$ we can pick $M_{0}>0$ such for any $w\in V_{d_{H}}$, if $\lVert\beta(T^{\prime}_{0})\varphi(T^{\prime}_{0}).w\rVert\geq M_{0}$, then
|$\left|\{t\in[T_{0}^{\prime},T]:\beta(t)\varphi(t).w\in\Psi\}\right|\leq\epsilon(T-T_{0}^{\prime}).$
@@
Since, $\lVert w_{l(T)}\rVert\to\infty$ as $T\to\infty$, we choose $T_{x}\geq T_{0}^{\prime}$, such that for all $T\geq T_{x}$, $\lVert\beta(T^{\prime}_{0})\varphi(T^{\prime}_{0}).w_{l(T)}\rVert\geq M_{0}$. Then, for any $T\geq T_{x}$, we have
(6.16) |$\left|\{t\in I(w_{l(T)}):\beta(t)\varphi(t).w_{l(T)}\in\Psi\}\right|\leq\epsilon(T-T_{0}^{\prime}).$
Thus, by combining (6.14), (6.15) and (6.16), for all $T\geq T_{x}$ we get
$\lvert J\cap[T_{0}^{\prime},T]\rvert\leq 2\epsilon(T-T_{0}^{\prime}).$
Therefore, (6.10) follows for $3\epsilon$ in place of $\epsilon$. ∎
6.2. Proving Theorem 1.8
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\mathcal{G}\leq G$ is a connected, closed Lie subgroup with a lattice $\Gamma\leq\mathcal{G}$. Suppose that $\varphi:[0,\infty)\to G$ be a continuous, unbounded, non-contracting curve for $G$ definable in a polynomially bounded o-minimal structure. Suppose that $\varphi\left([0,\infty)\right)\subset\mathcal{G}$. Let $H_{\varphi}$ be the hull of $\varphi$ in $G$ and let $\beta$ be a definable correcting curve in $G$, so that $\beta\varphi\subset H_{\varphi}$, see Definition 2.5. Then, by Proposition 5.13, $H_{\varphi}\subset\mathcal{G}$ and $\beta([0,\infty))\subset\mathcal{G}$. We choose an injective algebraic map $\iota:G\hookrightarrow\text{SL}(n,\mathbb{R})$ for some $n\geq 2$. Note that $\iota(\Gamma)$ is a discrete subgroup of $\text{SL}(n,\mathbb{R})$. Importantly, $\iota\circ\varphi$ is non-contracting in $\text{SL}(n,\mathbb{R})$ by Lemma 2.15, and $\iota(H_{\varphi})=H_{\iota\circ\varphi}$ by Lemma 2.8. By abuse of notations, in the following we replace $\varphi$ with $\iota\circ\varphi$, $\Gamma$ with $\iota(\Gamma)$ and $G$ with $\iota(G)$.
@@
Let $x_{0}\in\mathcal{G}/\Gamma\hookrightarrow\text{SL}(n,\mathbb{R})/\Gamma$, and fix a limiting measure $\mu$ of $\mu_{T,\beta\varphi,x_{0}}$ (as in (1.1)), with respect to the weak-$\ast$ topology on the space of probability measures on $\text{SL}(n,\mathbb{R})/\Gamma$, $T\to\infty$. We fix a sequence $\{T_{i}\}$ such that $T_{i}\to\infty$ and
(6.17) $\lim_{i\to\infty}\mu_{T_{i},\beta\varphi,x_{0}}=\mu.$
@@
By Theorem 4.5, $\mu$ is a probability measure. We let $W$ be the subgroup of $\text{SL}(n,\mathbb{R})$ generated by all unipotent one-parameter subgroups under which the limiting measure $\mu$ is invariant. By Proposition 5.1, $W$ has a positive dimension. Using Theorem 6.2, there exists $H\in\mathcal{H}$ (of smallest dimension) such that
@@
Define
(6.19) $N^{1}(H):=\eta_{H}^{-1}(p_{H})=\{g\in\text{SL}(n,\mathbb{R}):g.p_{H}=p_{H}\}.$
and recall
  • [Dani1993LimitDO, Theorem 3.4]: $N^{1}(H)\Gamma/\Gamma\subset X$ is closed.
@@
Fix $H\in\mathcal{H}$ such that (6.18) holds. Then there exists an $g_{0}\in\text{SL}(n,\mathbb{R})$ such that $x_{0}=\pi(g_{0})$ and
(6.20) $g_{0}^{-1}\beta(t)\varphi(t)g_{0}\subset N^{1}(H),~{}\forall t\geq 0,$
@@
In view of (6.1), due to (6.18), there is a compact subset $C\subset N(H,W)\smallsetminus S(H,W)$ such that $\mu(\pi(C))=\alpha$ for some $\alpha>0$, and $\pi(C)\cap\pi(S(H,W))=\varnothing$. We will now apply Theorem 6.4. We consider the compact subset $\text{K}_{1}\subset A_{H}$ given by $\text{K}_{1}:=C.p_{H}$ and fix $\epsilon:=\frac{\alpha}{2}$. Let $\mathcal{S}\subset\pi(S(H,W))$ be the closed subset given in Theorem 6.4. Now let $\Psi$ be an arbitrary open neighborhood of $\text{K}_{1}$, and let $K\subseteq X\smallsetminus\mathcal{S}$ be an arbitrary open neighborhood of $\pi(C)$, which exists as $\pi(C)$ is compact and $\mathcal{S}$ is closed. Then, $K\cap\pi(\eta_{H}^{-1}(\Psi))$ is an open neighborhood of $\pi(C)$. Since we assume (6.17), for all $i$ large enough we have
|$|\{t\in[0,T_{i}]:\beta(t)\varphi(t).x\in K\cap\pi(\eta_{H}^{-1}(\Psi))\}|\geq\epsilon T_{i}.$
@@
Namely, condition 6.4,(2) fails, and so the first outcome of Theorem 6.4,(1) must hold. That is, there is a $g_{0}\in\pi^{-1}(x)$ such that $g_{0}.p_{H}\in\Psi$ and
(6.22) $\beta(t)\varphi(t)g_{0}.p_{H}=g_{0}.p_{H},\text{ for }t\geq 0.$
@@
Namely, $\beta(t)\varphi(t)g_{0}\in g_{0}N^{1}(H)$ for all $t\geq T_{0}$. Since $\pi(N^{1}(H))$ is closed, we obtain that $\mu$ is supported in $\pi(g_{0}N^{1}(H))$.
Since $\pi^{-1}(x).p_{H}$ is discrete, and since we may choose the $\text{K}_{1}$-neighborhood $\Psi$ arbitrarily, we obtain that $g_{0}.p_{H}\in A_{H}$. Then, by Theorem 6.3,(1), we conclude that $g_{0}\in N(H,W)$. Then $g_{0}N^{1}(H)\subseteq N(H,W)$.
In particular, we conclude that $\mu(\pi(g_{0}N^{1}(H)\smallsetminus S(H,W))=1$. By [Mozes1995OnTS, Lemma 2.4], for each $g_{0}n\in g_{0}N^{1}(H)\smallsetminus S(H,W)$, the group
$H^{\prime}=(g_{0}n)H(g_{0}n)^{-1}=g_{0}Hg_{0}^{-1}$
is the smallest group containing $W$ such that the orbit $H^{\prime}\pi(g_{0}n)$ is closed. Then, by [Shah1991, Theorem 2.3], $W$ acts ergodically on $H^{\prime}\pi(g_{0}n)$ with respect to $g_{0}\mu_{H}$, where $\mu_{H}$ is the $H$ invariant measure on $H/H\cap\Gamma\cong H\Gamma/\Gamma$. Namely, $\mu$-almost every $W$-ergodic component of $\mu$ is $g_{0}Hg_{0}^{-1}$-invariant. By performing ergodic decomposition on $\mu$, we obtain that $\mu$ is $g_{0}Hg_{0}^{-1}$-invariant. Then, $\nu=g_{0}^{-1}\mu$ is $H$-invariant. So $H_{u}$ preserves $\nu$. Since $g_{0}^{-1}Wg_{0}\leq H$, we get $H_{u}=g_{0}^{-1}Wg_{0}$ by the maximality of $W$.
@@
We will use the following notations: $\psi(t):=g_{0}^{-1}\beta(t)\varphi(t)g_{0}$, for $t\geq 0$, $N:=N^{1}(H)$, $\Delta:=N^{1}(H)\cap\Gamma$, and $q:N\to N/H$ denotes the natural quotient map.
@@
Consider $\bar{N}:=q(N)$, $\bar{\Delta}:=q(\Delta)$. Since $H\Delta=\Delta H$ is closed, we have that $\bar{\Delta}\leq\bar{N}$ is discrete (since $\Delta H$ is closed, then $\Delta H/H\cong\Delta/\Delta\cap H$). We let $\bar{q}:N/\Delta\to\bar{N}/\bar{\Delta}$ be the natural map. Notice that the fibers of $\bar{q}$ are $H$-orbits. Namely,
$\bar{q}^{-1}(\bar{n}\bar{\Delta}))=nH\Delta/\Delta=Hn\Delta/\Delta.$
Since $\mu$ is a probability measure on $N/\Delta$, it follows that $\bar{\nu}:=\bar{q}_{*}\nu$ is a probability measure on $\bar{N}/\bar{\Delta}$.
@@
Suppose that $\psi(t)$ is unbounded in $N/H_{u}$. Then there exists an unipotent one-parameter subgroup $U\leq N$ such that $U$ is not a subgroup of $H$ and $\bar{\nu}$ is invariant under $q(U)$ and $q(U)$ is a (non-trivial) one-parameter subgroup.
@@
Thus, we may conclude the statement by applying Proposition 5.12, and Lemma 5.5 (see the proof of Lemma 5.6. The only difference here is that $N/H$ is not a subgroup of $\text{GL}(n,\mathbb{R})$). ∎
We now note the following, which is obtained by the ergodic decomposition for the $H$-action on $N/\Delta$ (see [EW11, Theorem 8.20])
Lemma 6.8.
Let $d(h\Delta)$ denote the volume element of the $H$-invariant probability measure on $H\Delta/\Delta$. For $f\in C_{c}(N/\Delta)$ let:
(6.23) $P(f)(q(n)\bar{\Delta}):=\int_{H\Delta/\Delta}f(nh\Delta)d(h\Delta).$
Then,
@@
We now have the following key observation.
@@
It holds that $\psi(t)\in H_{u}$ for all $t\geq T_{0}$, and $\nu$ is the $H$-invariant probability on $H\pi(e)$, where $\pi(e)$ is the identity coset in $X=\text{SL}(n,\mathbb{R})/\Gamma$. Moreover, $H$ is the smallest subgroup such that $g_{0}^{-1}H_{\varphi}g_{0}\leq H$ and such that the orbit $H\pi(e)$ is closed.
@@
We first show that $\psi(t)$ is bounded in $N/H_{u}$. In fact, if $\psi(t)$ is unbounded in $N/H_{u}$, then by Lemma 6.7, $\bar{\nu}$ is invariant under $q(U)$, where $U$ is a one-parameter unipotent subgroup and $q(U)$ is a non-trivial one-parameter subgroup of $\bar{N}=N/H$. It then follows from Lemma 6.8 than $\nu$ is invariant by $U$ and $U$ is not a subgroup of $H$. Since $\nu$ was invariant by $H$ as well, we get that $\nu$ is invariant by a group generated by one-parameter subgroups which strictly includes $H_{u}$. This is a contradiction since $H_{u}$ is a maximal group with this property.
Since the image of $\psi(t)=g_{0}^{-1}\beta(t)\varphi(t)g_{0}$ is boundned in $\text{SL}(n,\mathbb{R})/H_{u}$, we get that the image of $\beta(t)\varphi(t)$ is bounded in $\text{SL}(n,\mathbb{R})/g_{0}H_{u}g_{0}^{-1}$. Thus, since $\text{zcl}(H_{u})$ is observable (see [Observable_group_grosshans, Corollary 2.8]), we conclude by Theorem 2.4 that $H_{\varphi}\leq g_{0}H_{u}g_{0}^{-1}$, and in particular $\beta(t)\varphi(t)\in g_{0}H_{u}g_{0}^{-1}$ for all $t$. Namely, $\psi(t)\in g_{0}^{-1}H_{\varphi}g_{0}\leq H_{u}$ for all $t\geq T_{0}$. In particular, $\bar{\nu}$ is the Dirac measure at the identity coset, and so $\nu=\mu_{H}$ by Lemma 6.8. In particular $\{\psi(t)\pi(e)\}_{t\geq 0}$ is dense in $H\pi(e)$. In particular $g_{0}^{-1}H_{\varphi}g_{0}\pi(e)$ is dense in $H\pi(e)$, which finishes the proof.
The proof of the main Theorem 1.8 follows directly from the above corollary.

### v2

@@ 6.2. O-minimal curves in representations – a dichotomy theorem
By Proposition 1.13, and recalling (1.6), given $\epsilon>0$ we can pick $M_{0}>0$ such for any $w\in V_{H}$, if $\lVert\psi(T^{\prime}_{0})w\rVert\geq M_{0}$, then
(6.16) |$\left|\{s\in[T_{0}^{\prime},T]:\psi(s)w\in\Psi\}\right|\leq\epsilon(T-T_{0}^{\prime}).$
@@
Due to the above claim, let $i_{x}\in\mathbb{N}$ be such that $\lVert w_{t_{i}}\rVert\geq\lVert\psi(T_{0}^{\prime})^{-1}\rVert\cdot M_{0}$ for all $i\geq i_{x}$. Set $T_{x}=t_{i_{x}}$. Let $T\geq T_{x}$. Since $t_{j}\to\infty$ as $j\to\infty$, $i:=\max\{j:t_{j}\leq T\}<\infty$. Then $i\geq i_{x}$, so $\lVert\psi(T^{\prime}_{0})w_{t_{i}}\rVert\geq M_{0}$. By (6.16),
(6.17) $\lvert J\cap[t_{i},T]\rvert\leq\lvert\{s\in[t_{i},T]:\psi(s)w_{t_{i}}\in\Psi\}\rvert\leq\epsilon(T-T_{0}^{\prime}).$
By combining (6.15) and (6.17), we obtain (6.11). ∎
6.3. Proof of Theorem 1.7
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\mathcal{G}\subseteq G$ is a connected, closed Lie subgroup with a lattice $\Gamma\subseteq\mathcal{G}$. Let $\varphi:[0,\infty)\to G$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve for $G$ such that $\varphi\left([0,\infty)\right)\subseteq\mathcal{G}$. Let $H_{\varphi}$ be the hull of $\varphi$ in $G$ and let $\beta$ be a definable correcting curve in $G$, so that $\beta\varphi\subseteq H_{\varphi}$, see Definition 2.6. Then, by Proposition 5.13, $H_{\varphi}\subseteq\mathcal{G}$ and $\beta([0,\infty))\subseteq\mathcal{G}$.
We choose an injective algebraic map $\iota:G\hookrightarrow\text{SL}(n,\mathbb{R})$ for some $n\geq 2$. Note that $\iota(\Gamma)$ is a discrete subgroup of $\text{SL}(n,\mathbb{R})$. Importantly, $\iota\circ\varphi$ is non-contracting in $\text{SL}(n,\mathbb{R})$ by Lemma 2.17, and $\iota(H_{\varphi})=H_{\iota\circ\varphi}$ by Lemma 2.10. By abuse of notations, in the following we replace $\varphi$ with $\iota\circ\varphi$, replace $\Gamma$ with $\iota(\Gamma)$ and replace $G$ with $\iota(G)$.
@@
Let $x_{0}\in\mathcal{G}/\Gamma\hookrightarrow\text{SL}(n,\mathbb{R})/\Gamma$, and fix a limiting measure $\mu$ as $T\to\infty$ of $\mu_{T,\beta\varphi,x_{0}}$ (defined in (1.1)), with respect to the weak-$\ast$ topology on the space of probability measures on $\text{SL}(n,\mathbb{R})/\Gamma$. We fix a sequence $\{T_{i}\}$ such that $T_{i}\to\infty$ and
[…]
By Theorem 4.5, $\mu$ is a probability measure.
@@
Let $W$ be the subgroup of $\text{SL}(n,\mathbb{R})$ generated by all unipotent one-parameter subgroups under which the limiting measure $\mu$ is invariant. By Proposition 5.1, $\dim(W)>0$. Using Theorem 6.3, there exists $H\in\mathcal{H}$ such that
@@
Here $H\in\mathcal{H}$ is of smallest dimension such that $\mu(\pi(N(H,W)))>0$.
In view of (6.7), we define
(6.20) $N^{1}(H)$ $:=\eta_{H}^{-1}(p_{H})=\{g\in\text{SL}(n,\mathbb{R}):g\cdot p_{H}=p_{H}\}$
=$=\{g\in N(H):\det((\text{Ad}g)|_{\mathfrak{h}})=1\},$
where $p_{H}$ and $\eta_{H}$ are as in (6.3). Since $H\cap\Gamma$ is a lattice in $H$, $H$ is unimodular (see [RA72, Remark 1.9]). So, $H\subseteq N^{1}(H)$.
The remainder of the proof proceeds as follows. In Lemma 6.7, we show that there exists an element $g_{0}\in\text{SL}(n,\mathbb{R})$ such that $g_{0}\Gamma=x_{0}$ and
$\{\beta(t)\varphi(t)g_{0}\}_{t\geq 0}\subseteq g_{0}N^{1}(H).$
We first note that $\Gamma N^{1}(H)=\eta_{H}^{-1}(\Gamma\cdot p_{H})$ is closed, as $\Gamma\cdot p_{H}$ is discrete in $V_{H}$ (see Theorem 6.4, (2)). Consequently, the quotient $N^{1}(H)\Gamma/\Gamma\subseteq X$ is also closed.
By (6.18), it follows that $\mu$ is supported on $g_{0}N^{1}(H)x_{0}$. One may view $g_{0}N^{1}(H)\Gamma/\Gamma$ as foliated by orbits of $g_{0}Hg_{0}^{-1}$, and we will show that $\mu$ is invariant under $g_{0}Hg_{0}^{-1}$. The key step is, to prove that $\mu$ is actually supported on the single orbit of $g_{0}Hg_{0}^{-1}$ passing through $x_{0}$, a fact established in Corollary 6.9. This will then immediately imply our main result, Theorem 1.7, as it shows that
$\overline{H_{\varphi}x_{0}}=g_{0}Hg_{0}^{-1}x_{0},$
and that $\beta\varphi$ is equidistributed on this orbit.
@@
Fix $H\in\mathcal{H}$ such that (6.19) holds. Then, there exists $g_{0}\in\text{SL}(n,\mathbb{R})$ such that $x_{0}=\pi(g_{0}):=g_{0}\Gamma$ and
(6.21) $\left\{g_{0}^{-1}\beta(t)\varphi(t)g_{0}\right\}_{t\geq 0}\subseteq N^{1}(H).$
@@
In view of (6.1), due to (6.19), there is a compact subset $C\subseteq N(H,W)\setminus S(H,W)$ such that $\mu(\pi(C))=\alpha$ for some $\alpha>0$, and $\pi(C)\cap\pi(S(H,W))=\varnothing$. Consider the compact subset $\text{K}_{1}\subseteq A_{H}$ given by $\text{K}_{1}:=C\cdot p_{H}$. We will now apply Theorem 6.5 to $x_{0}\in X$.
We show that outcome (2) of Theorem 6.5 cannot hold. Let $\epsilon:=\frac{\alpha}{2}$. Denote by $\mathcal{S}\subseteq\pi(S(H,W))$ the closed subset given by outcome (2) of Theorem 6.5. Let $K\subseteq X\setminus\mathcal{S}$ be an open neighborhood of $\pi(C)$, which exists as $\pi(C)$ is compact and $\mathcal{S}$ is closed. Also, fix an arbitrary compact neighborhood $U$ of $\text{K}_{1}$ in $V_{H}$. Then, $K\cap\pi(\eta_{H}^{-1}(U))$ is an open neighborhood of $\pi(C)$. By (6.18), for all $i$ large enough,
|$|\{t\in[0,T_{i}]:\beta(t)\varphi(t)x_{0}\in K\cap\pi(\eta_{H}^{-1}(U))\}|\geq\epsilon T_{i},$
which contradicts outcome (2) of Theorem 6.5.
@@
Therefore, the first outcome (1) of Theorem 6.5 must hold. That is, we can pick $g_{0}\in\pi^{-1}(x_{0})$ such that $g_{0}\cdot p_{H}\in\text{K}_{1}\subseteq A_{H}$ and such that:
(6.23) $\beta(t)\varphi(t)g_{0}\cdot p_{H}=g_{0}\cdot p_{H},\text{ for }t\geq 0.$
@@
Namely, $\beta(t)\varphi(t)g_{0}\in g_{0}N^{1}(H)$ for all $t\geq 0$. Since $\pi(N^{1}(H))$ is closed, we obtain that $\mu$ is supported on a subset of $\pi(g_{0}N^{1}(H))$. In fact, we claim that $\mu(\pi(g_{0}N^{1}(H)))=1$
To see this, observe that since $g_{0}\cdot p_{H}\in A_{H}$, then $g_{0}\in N(H,W)$ by Theorem 6.4,(1). Moreover, $g_{0}N^{1}(H)\subseteq N(H,W)$, since $N(H,W)$ is right-$N^{1}(H)$ invariant.
In particular, we conclude that $\mu(\pi(g_{0}N^{1}(H)\setminus S(H,W)))=1$ (see (6.19)). Therefore, by [Mozes1995OnTS, Lemma 2.4], almost every $W$-ergodic component of $\mu$ is invariant under $gHg^{-1}=g_{0}Hg_{0}^{-1}$ for some $g\in g_{0}N^{1}(H)\setminus S(H,W)$. Therefore $\mu$ is $g_{0}Hg_{0}^{-1}$-invariant. Then, $\nu=(g_{0}^{-1})_{*}\mu$ is $H$-invariant, and especially, $H_{u}$ preserves $\nu$.
Finally, we claim that $H_{u}=g_{0}^{-1}Wg_{0}\subseteq H$. To see this equality, recall that $g_{0}\in N(H,W)$, so that $g_{0}^{-1}Wg_{0}\subseteq H$, and recall that $W$ is chosen to be the subgroup of $\text{SL}(n,\mathbb{R})$ generated by all unipotent one-parameter subgroups preserving $\mu$. ∎
@@
We will use the following notations: $\psi(t):=g_{0}^{-1}\beta(t)\varphi(t)g_{0}$, for $t\geq 0$, $N:=N^{1}(H)$, $\Delta:=N^{1}(H)\cap\Gamma$ and $q:N\to N/H$ denotes the natural quotient map.
@@
Consider $\bar{N}:=q(N)$, $\bar{\Delta}:=q(\Delta)$. Since $H\Delta=\Delta H$ is closed, we have that $\bar{\Delta}\leq\bar{N}$ is discrete (since $\Delta H$ is closed, then $q(\Delta)=\Delta H/H\cong\Delta/\Delta\cap H$). We let $\bar{q}:N/\Delta\to\bar{N}/\bar{\Delta}$ be the natural map. Notice that the fibers of $\bar{q}$ are $H$-orbits. Namely,
$\bar{q}^{-1}(\bar{n}\bar{\Delta})=nH\Delta/\Delta=Hn\Delta/\Delta.$
Since $\nu$ is a probability measure on $N/\Delta\cong\pi(N^{1}(H))$, it follows that $\bar{\nu}:=\bar{q}_{*}\nu$ is a probability measure on $\bar{N}/\bar{\Delta}$.
@@
Suppose that $\psi(t)$ is unbounded in $N/H_{u}$. Then there exists a unipotent one-parameter subgroup $U\leq N$ such that $U$ is not contained in $H$, the measure $\bar{\nu}$ is invariant under $q(U)$, and $q(U)$ is a nontrivial one-parameter subgroup.
Moreover, $\nu$ is $U$-invartiant.
@@
Thus, we may conclude the statement by applying Proposition 5.12, and Lemma 5.5 (see the proof of Lemma 5.6. The only difference here is that $N/H$ is not a subgroup of $\text{GL}(n,\mathbb{R})$).
By the ergodic decomposition for the $H$-action on $N/\Delta$ (see e.g. [EW11, Theorem 8.20]), for all $f\in C_{c}(N/\Delta)$,
@@
where $\mu_{H\Delta/\Delta}$ denotes the $H$-invariant probability measure on $H\Delta/\Delta$.
Therefore, since $\bar{\nu}$ is $q(U)$-invariant, $\nu$ is $U$-invariant. ∎
@@
It holds that $\psi(t)\in H_{u}$ for all $t\geq 0$, and $\nu$ is the $H$-invariant probability on $H\Delta/\Delta\cong H\Gamma/\Gamma$. Moreover, $H$ is the smallest subgroup such that $g_{0}^{-1}H_{\varphi}g_{0}\subseteq H$ and that the orbit $H\Delta/\Delta$ is closed.
@@
We first show that $\psi(t)$ is bounded in $N/H_{u}$. Suppose $\psi(t)$ is unbounded in $N/H_{u}$. Then by Lemma 6.8, $\nu$ is invariant by the group generated by $U$ and $H$, which strictly includes $H_{u}$. This is a contradiction, because $H_{u}$ is the maximal subgroup generated by unipotent one-parameter subgroups preserving $\nu$.
Now, since the image of $\psi(t)=g_{0}^{-1}\beta(t)\varphi(t)g_{0}$ is boundned in $\text{SL}(n,\mathbb{R})/H_{u}$, we get that the image of $\beta(t)\varphi(t)$ is bounded in $\text{SL}(n,\mathbb{R})/g_{0}H_{u}g_{0}^{-1}$. Thus, since $\text{zcl}(H_{u})$ is observable (see Lemma 2.3), by minimality of the hull (see Theorem 2.5), $H_{\varphi}\subseteq g_{0}H_{u}g_{0}^{-1}$, and in particular $\beta(t)\varphi(t)\in g_{0}H_{u}g_{0}^{-1}$ for all $t$. Namely, $\psi(t)\in g_{0}^{-1}H_{\varphi}g_{0}\subseteq H_{u}$ for all $t\geq 0$. In particular, $\bar{\nu}$ is the Dirac measure at the identity coset. Therefore $\nu=\mu_{H\Delta/\Delta}$ by (6.24). In particular $\{\psi(t)\Delta\}_{t\geq 0}$ is dense in $H\Delta/\Delta$. In particular $g_{0}^{-1}H_{\varphi}g_{0}\Delta$ is dense in $H\Delta/\Delta$, which finishes the proof.
Theorem 1.7 follows directly from the above corollary.